# A Benchmark for Multi-modal Foundation Models on Low-level Vision: from Single Images to Pairs

### v1

@@ III Experiment
In Q-Bench, we evaluate the performance of up to 20 up-to-date popular and competitive open-source as well as 4 close-source commercial MLLMs under zero-shot settings.
@@ III-A Findings on Perception
For a holistic examination of the perception ability of MLLMs, we evaluate the multi-choice correctness of MLLMs on different sub-categories of the LLVision dataset, which is equally divided as dev (will be released) and test (will keep private) subsets as shown in Table II and Table III respectively. Only the MLLMs that support multiple images input are included for the perception-pair ability benchmark.
@@ III-A1 Perception for Single Images
a) We are glad that the majority of MLLMs can significantly outperform random guess on all sub-categories as shown in Table II. Considering that all participating MLLMs are without any explicit training on low-level visual attributes, these results show strong potentials for these general-purpose models when further fine-tuned with respective low-level datasets. b) Among all open-source MLLMs, the recently-released SPHINX reaches the best accuracy on this question-answering task, followed by Emu2-Chat and InternLM-XComposer-VL, which show rather close results. By achieving more than 64% accuracy on both subsets, these models show exciting potential as robust low-level visual assistants in the future. c) Another key observation is that almost all methods perceive worse on distortions than other low-level attributes, which indicates that distortion questions are relatively more challenging. d) Close-source MLLMs and Humans. It is widely acknowledged that commercial close-source MLLMs are the leading models in various tasks. To evaluate the low-level perception abilities of these MLLMs, we gauge the accuracy of Qwen-VL-Plus (Alibaba), Qwen-VL-Max (Alibaba), Gemini-Pro (Google), and GPT-4V (OpenAI) on the subsets of LLVision dataset. All close-source MLLMs achieve superior performance than all open-source MLLMs on the test subset, which indicates that open-source MLLMs still fall behind on low-level visual ability. GPT-4V exhibits the most competitive performance and outperforms the best open-source MLLM (SPHINX) by a large margin (+6%), and on par accuracy with the Junior-level Human. Despite its prowess, there is still a way to go for GPT-4V before it can match the overall proficiency of the Senior-level Human (with experiences on low-level visual tasks, 7% better than GPT-4V). Furthermore, across all categories, the results show that GPT-4V, much like its open-source counterparts, faces challenges in recognizing distortions.
@@
| InfiMM (Zephyr-7B) [56] | 29.61% | 62.32% | 7.77% | 0.77 | 29.25% | 31.90% | 38.51% | 1.08 | 2.16% | 22.72% | 74.58% | 1.71 | 3.58 |
| Emu2-Chat (LLaMA-33B) [25] | 20.01% | 52.77% | 27.22% | 1.07 | 24.66% | 27.12% | 48.22% | 1.24 | 1.21% | 9.91% | 88.88% | 1.88 | 4.19 |
| Fuyu-8B (Persimmon-8B) [57] | 25.54% | 61.00% | 13.46% | 0.88 | 41.96% | 32.76% | 25.28% | 0.83 | 2.99% | 11.34% | 85.67% | 1.82 | 3.53 |
| BakLLava (Mistral-7B) [26] | 24.31% | 51.22% | 24.47% | 1.00 | 49.23% | 24.11% | 26.66% | 0.77 | 1.25% | 36.22% | 62.53% | 1.61 | 3.38 |
| SPHINX [58] | 27.96% | 64.36% | 7.33% | 0.79 | 26.16% | 32.42% | 41.01% | 1.14 | 1.69% | 23.00% | 74.61% | 1.72 | 3.65 |
| mPLUG-Owl2 (LLaMA-7B) [59] | 27.71% | 38.58% | 33.71% | 1.06 | 28.11% | 19.78% | 52.11% | 1.24 | 7.91% | 48.18% | 43.91% | 1.36 | 3.67 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [60] | 27.48% | 54.74% | 17.78% | 0.90 | 30.51% | 26.04% | 43.45% | 1.13 | 10.85% | 60.34% | 28.81% | 1.18 | 3.21 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [60] | 27.68% | 53.78% | 18.55% | 0.91 | 25.45% | 21.47% | 53.08% | 1.28 | 6.31% | 58.75% | 34.94% | 1.29 | 3.47 |
| InternLM-XComposer-VL (InternLM) [61] | 19.94% | 51.82% | 28.24% | 1.08 | 22.59% | 28.99% | 48.42% | 1.26 | 1.05% | 10.62% | 88.32% | 1.87 | 4.21 |
| IDEFICS-Instruct (LLaMA-7B) [62] | 28.91% | 59.16% | 11.93% | 0.83 | 34.68% | 27.86% | 37.46% | 1.03 | 3.90% | 59.66% | 36.44% | 1.33 | 3.18 |
| Qwen-VL (QwenLM) [63] | 26.34% | 49.13% | 24.53% | 0.98 | 50.62% | 23.44% | 25.94% | 0.75 | 0.73% | 35.56% | 63.72% | 1.63 | 3.36 |
| Shikra (Vicuna-7B) [64] | 21.14% | 68.33% | 10.52% | 0.89 | 30.33% | 28.30% | 41.37% | 1.11 | 1.14% | 64.36% | 34.50% | 1.33 | 3.34 |
| Otter-v1 (MPT-7B) [6] | 22.38% | 59.36% | 18.25% | 0.96 | 40.68% | 35.99% | 23.33% | 0.83 | 1.95% | 13.20% | 84.85% | 1.83 | 3.61 |
| Kosmos-2 [9] | 8.76% | 70.91% | 20.33% | 1.12 | 29.45% | 34.75% | 35.81% | 1.06 | 0.16% | 14.77% | 85.06% | 1.85 | 4.03 |
| InstructBLIP (Flan-T5-XL) [5] | 23.16% | 66.44% | 10.40% | 0.87 | 34.85% | 26.03% | 39.12% | 1.04 | 14.71% | 59.87% | 25.42% | 1.11 | 3.02 |
| InstructBLIP (Vicuna-7B) [5] | 29.73% | 61.47% | 8.80% | 0.79 | 27.84% | 23.52% | 48.65% | 1.21 | 27.40% | 61.29% | 11.31% | 0.84 | 2.84 |
| VisualGLM-6B (GLM-6B) [65] | 30.75% | 56.64% | 12.61% | 0.82 | 38.64% | 26.18% | 35.18% | 0.97 | 6.14% | 67.15% | 26.71% | 1.21 | 2.99 |
| mPLUG-Owl (LLaMA-7B) [52] | 28.28% | 37.69% | 34.03% | 1.06 | 26.75% | 18.18% | 55.07% | 1.28 | 3.03% | 33.82% | 63.15% | 1.60 | 3.94 |
| LLaMA-Adapter-V2 [66] | 30.44% | 53.99% | 15.57% | 0.85 | 29.41% | 25.79% | 44.80% | 1.15 | 1.50% | 52.75% | 45.75% | 1.44 | 3.45 |
| LLaVA-v1 (Vicuna-13B) [3] | 34.10% | 40.52% | 25.39% | 0.91 | 30.02% | 15.15% | 54.83% | 1.25 | 1.06% | 38.03% | 60.91% | 1.60 | 3.76 |
| MiniGPT-4 (Vicuna-13B) [4] | 34.01% | 32.15% | 33.85% | 1.00 | 29.20% | 15.27% | 55.53% | 1.26 | 6.88% | 45.65% | 47.48% | 1.41 | 3.67 |
@@
| InfiMM (Zephyr-7B) [56] | 30.75% | 62.66% | 6.22% | 0.75 | 34.17% | 38.84% | 26.35% | 0.91 | 2.57% | 30.84% | 65.28% | 1.61 | 3.28 |
| Emu2-Chat (LLaMA-33B) [25] | 41.25% | 54.33% | 4.42% | 0.63 | 38.11% | 36.41% | 25.48% | 0.87 | 4.12% | 38.61% | 57.27% | 1.53 | 3.03 |
| Fuyu-8B (Persimmon-8B) [57] | 37.95% | 52.17% | 9.11% | 0.70 | 37.68% | 37.33% | 23.73% | 0.84 | 3.95% | 31.15% | 62.84% | 1.56 | 3.12 |
| BakLLava (Mistral-7B) [26] | 29.46% | 59.77% | 10.57% | 0.80 | 40.0% | 38.08% | 21.33% | 0.80 | 2.26% | 15.06% | 82.04% | 1.79 | 3.40 |
| mPLUG-Owl2 (Q-Instruct) [59] | 15.25% | 65.76% | 18.32% | 1.02 | 39.44% | 40.18% | 19.62% | 0.79 | 0.09% | 9.86% | 89.02% | 1.87 | 3.69 |
| mPLUG-Owl2 (LLaMA-7B) [59] | 19.43% | 65.54% | 14.45% | 0.94 | 30.94% | 43.71% | 24.63% | 0.92 | 3.79% | 26.94% | 68.28% | 1.63 | 3.50 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [60] | 19.68% | 72.57% | 7.19% | 0.86 | 38.00% | 40.04% | 20.97% | 0.82 | 2.13% | 39.77% | 56.66% | 1.53 | 3.22 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [60] | 18.77% | 73.44% | 7.79% | 0.89 | 34.66% | 38.72% | 26.62% | 0.92 | 1.02% | 34.59% | 64.39% | 1.63 | 3.44 |
@@ III-A2 Perception for Image Pairs
Perception for image pairs is far more difficult for MLLMs since this task not only requires MLLMs to have stable low-level visual capabilities, but also requires MLLMs to be able to analyze two images simultaneously and conduct discerning comparisons. To enrich the MLLM diversity, we further include the mPLUG-Owl2 fine-tuned with the single image low-level visual dataset Q-Instruct [67] for comparison. The performance is exhibited in Table III. With closer inspections, we can obtain several interesting findings. a) Open-source MLLMs are poor low-level comparators. It seems that although they might show strong performance for single image perception, they are quite confused by the image pairs. Most of them get worse performance on the Compare subset than the Joint subset, which further confirms this point. For mPLUG-Owl2 (Q-Instruct), despite being fine-tuned with the single image low-level visual dataset Q-Instruct [67], the overall performance improvement from the low-level knowledge infusion of single images is relatively weak. This also suggests that there is a necessity to build open-source low-level datasets for multiple images to cultivate the corresponding capabilities of open-source MLLMs. b) Close-source MLLMs are more robust in this task. This may be because these close-source MLLMs are supported by training on multiple-image data, allowing them to make better comparative judgments. Particularly with GPT-4V, its performance in the compare subset is significantly higher than in the joint subset, and it far exceeds all other models, even reaching the level of a junior human. c) Perception for image pairs is easier for humans. Comparing image pairs is simpler for humans, as the answers to related questions tend to be more objective. Especially for junior-level humans with no professional experience, they may have stronger subjectivity in grasping absolute sensations, but it is easier to remain objective when they are faced with comparison concepts. For example, it’s difficult for a junior-level human to judge whether the lighting in a dimly lit single image is appropriate. However, if presented with another image with even weaker lighting, they can easily determine which image is worse. This may explain the notable 6% improvements for junior-level human from single images to image pairs.
@@ III-B1 Description for single images
| NIQE [68] | 0.316/0.377 | 0.693/0.669 | 0.211/0.288 | 0.480/0.451 | 0.075/0.056 | 0.562/0.517 | 0.374/0.428 | 0.387/0.398 |
| CLIP-ViT-Large-14 [69] | 0.468/0.505 | 0.385/0.389 | 0.218/0.237 | 0.307/0.308 | 0.285/0.290 | 0.436/0.458 | 0.376/0.388 | 0.354/0.368 |
@@
| Emu2-Chat (LLaMA-33B) [25] | 0.664/0.714 | 0.712/0.698 | 0.355/0.341 | 0.597/0.611 | 0.224/0.269 | 0.759/0.751 | 0.841/0.790 | 0.593/0.596 |
| Fuyu-8B (Persimmon-8B) [57] | 0.124/0.123 | 0.125/0.179 | 0.164/0.133 | 0.225/0.176 | 0.118/0.116 | 0.368/0.317 | 0.099/0.088 | 0.174/0.161 |
| BakLLava (Mistral-7B) [26] | 0.389/0.390 | 0.406/0.398 | 0.227/0.216 | 0.335/0.337 | 0.179/0.209 | 0.542/0.561 | 0.344/0.361 | 0.346/0.353 |
| mPLUG-Owl2 (LLaMA-7B) [59] | 0.196/0.252 | 0.589/0.614 | 0.217/0.286 | 0.293/0.342 | -0.024/-0.032 | 0.473/0.492 | 0.541/0.546 | 0.326/0.357 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [60] | 0.463/0.459 | 0.443/0.467 | 0.305/0.321 | 0.344/0.358 | 0.321/0.333 | 0.672/0.738 | 0.417/0.440 | 0.424/0.445 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [60] | 0.448/0.460 | 0.563/0.584 | 0.310/0.339 | 0.445/0.481 | 0.285/0.297 | 0.664/0.754 | 0.390/0.400 | 0.444/0.474 |
| InternLM-XComposer-VL (InternLM) [61] | 0.564/0.615 | 0.730/0.750 | 0.360/0.416 | 0.612/0.676 | 0.243/0.265 | 0.732/0.775 | 0.546/0.572 | 0.541/0.581 |
| IDEFICS-Instruct (LLaMA-7B) [62] | 0.375/0.400 | 0.474/0.484 | 0.235/0.240 | 0.409/0.428 | 0.244/0.227 | 0.562/0.622 | 0.370/0.373 | 0.381/0.396 |
| Qwen-VL (QwenLM) [63] | 0.470/0.546 | 0.676/0.669 | 0.298/0.338 | 0.504/0.532 | 0.273/0.284 | 0.617/0.686 | 0.486/0.486 | 0.475/0.506 |
| Shikra (Vicuna-7B) [64] | 0.314/0.307 | 0.320/0.337 | 0.237/0.241 | 0.322/0.336 | 0.198/0.201 | 0.640/0.661 | 0.324/0.332 | 0.336/0.345 |
| Otter-v1 (MPT-7B) [6] | 0.406/0.406 | 0.436/0.441 | 0.143/0.142 | -0.008/0.018 | 0.254/0.264 | 0.475/0.481 | 0.557/0.577 | 0.323/0.333 |
| Kosmos-2 [9] | 0.255/0.281 | 0.644/0.641 | 0.196/0.195 | 0.358/0.368 | 0.210/0.225 | 0.489/0.491 | 0.359/0.365 | 0.359/0.367 |
| InstructBLIP (Flan-T5-XL) [5] | 0.334/0.362 | 0.582/0.599 | 0.248/0.267 | 0.113/0.113 | 0.167/0.188 | 0.378/0.400 | 0.211/0.179 | 0.290/0.301 |
| InstructBLIP (Vicuna-7B) [5] | 0.359/0.437 | 0.683/0.689 | 0.200/0.283 | 0.253/0.367 | 0.263/0.304 | 0.629/0.663 | 0.337/0.382 | 0.389/0.446 |
| VisualGLM-6B (GLM-6B) [65] | 0.247/0.234 | 0.498/0.507 | 0.146/0.154 | 0.110/0.116 | 0.209/0.183 | 0.342/0.349 | 0.127/0.131 | 0.240/0.239 |
| mPLUG-Owl (LLaMA-7B) [52] | 0.409/0.427 | 0.634/0.644 | 0.241/0.271 | 0.437/0.487 | 0.148/0.180 | 0.687/0.711 | 0.466/0.486 | 0.432/0.458 |
| LLaMA-Adapter-V2 [66] | 0.354/0.363 | 0.464/0.506 | 0.275/0.329 | 0.298/0.360 | 0.257/0.271 | 0.604/0.666 | 0.412/0.425 | 0.381/0.417 |
| LLaVA-v1 (Vicuna-13B) [3] | 0.462/0.457 | 0.442/0.462 | 0.264/0.280 | 0.404/0.417 | 0.208/0.237 | 0.626/0.684 | 0.349/0.372 | 0.394/0.416 |
| MiniGPT-4 (Vicuna-13B) [4] | 0.239/0.257 | 0.238/0.253 | 0.170/0.183 | 0.339/0.340 | 0.252/0.246 | 0.572/0.591 | 0.239/0.233 | 0.293/0.300 |

### v2

@@ III Experiment
In Q-Bench+, we evaluate the performance of up to 20 up-to-date popular and competitive open-source as well as 4 close-source commercial MLLMs under zero-shot settings.
@@ III-A Findings on Perception
For a holistic examination of the perception ability of MLLMs, we evaluate the multi-choice correctness of MLLMs on different sub-categories of the LLVision+ dataset, which is equally divided as dev (will be released) and test (will keep private) subsets as shown in Table II and Table III respectively. Only the MLLMs that support multiple images input are included for the perception-pair ability benchmark.
@@ III-A1 Perception for Single Images
a) We are glad that the majority of MLLMs can significantly outperform random guess on all sub-categories as shown in Table II. Considering that all participating MLLMs are without any explicit training on low-level visual attributes, these results show strong potentials for these general-purpose models when further fine-tuned with respective low-level datasets. b) Among all open-source MLLMs, the recently-released SPHINX reaches the best accuracy on this question-answering task, followed by Emu2-Chat and InternLM-XComposer-VL, which show rather close results. By achieving more than 64% accuracy on both subsets, these models show exciting potential as robust low-level visual assistants in the future. c) Another key observation is that almost all methods perceive worse on distortions than other low-level attributes, which indicates that distortion questions are relatively more challenging. d) Close-source MLLMs and Humans. It is widely acknowledged that commercial close-source MLLMs are the leading models in various tasks. To evaluate the low-level perception abilities of these MLLMs, we gauge the accuracy of Qwen-VL-Plus (Alibaba), Qwen-VL-Max (Alibaba), Gemini-Pro (Google), and GPT-4V (OpenAI) on the subsets of LLVision+ dataset. All close-source MLLMs achieve superior performance than all open-source MLLMs on the test subset, which indicates that open-source MLLMs still fall behind on low-level visual ability. GPT-4V exhibits the most competitive performance and outperforms the best open-source MLLM (SPHINX) by a large margin (+6%), and on par accuracy with the Junior-level Human. Despite its prowess, there is still a way to go for GPT-4V before it can match the overall proficiency of the Senior-level Human (with experiences on low-level visual tasks, 7% better than GPT-4V). Furthermore, across all categories, the results show that GPT-4V, much like its open-source counterparts, faces challenges in recognizing distortions.
@@
| InfiMM (Zephyr-7B) [InfiMM] | 29.61% | 62.32% | 7.77% | 0.77 | 29.25% | 31.90% | 38.51% | 1.08 | 2.16% | 22.72% | 74.58% | 1.71 | 3.58 |
| Emu2-Chat (LLaMA-33B) [emu2] | 20.01% | 52.77% | 27.22% | 1.07 | 24.66% | 27.12% | 48.22% | 1.24 | 1.21% | 9.91% | 88.88% | 1.88 | 4.19 |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 25.54% | 61.00% | 13.46% | 0.88 | 41.96% | 32.76% | 25.28% | 0.83 | 2.99% | 11.34% | 85.67% | 1.82 | 3.53 |
| BakLLava (Mistral-7B) [bakllava] | 24.31% | 51.22% | 24.47% | 1.00 | 49.23% | 24.11% | 26.66% | 0.77 | 1.25% | 36.22% | 62.53% | 1.61 | 3.38 |
| SPHINX [sphinx] | 27.96% | 64.36% | 7.33% | 0.79 | 26.16% | 32.42% | 41.01% | 1.14 | 1.69% | 23.00% | 74.61% | 1.72 | 3.65 |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 27.71% | 38.58% | 33.71% | 1.06 | 28.11% | 19.78% | 52.11% | 1.24 | 7.91% | 48.18% | 43.91% | 1.36 | 3.67 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 27.48% | 54.74% | 17.78% | 0.90 | 30.51% | 26.04% | 43.45% | 1.13 | 10.85% | 60.34% | 28.81% | 1.18 | 3.21 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 27.68% | 53.78% | 18.55% | 0.91 | 25.45% | 21.47% | 53.08% | 1.28 | 6.31% | 58.75% | 34.94% | 1.29 | 3.47 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 19.94% | 51.82% | 28.24% | 1.08 | 22.59% | 28.99% | 48.42% | 1.26 | 1.05% | 10.62% | 88.32% | 1.87 | 4.21 |
| IDEFICS-Instruct (LLaMA-7B) [idefics] | 28.91% | 59.16% | 11.93% | 0.83 | 34.68% | 27.86% | 37.46% | 1.03 | 3.90% | 59.66% | 36.44% | 1.33 | 3.18 |
| Qwen-VL (QwenLM) [Qwen-VL] | 26.34% | 49.13% | 24.53% | 0.98 | 50.62% | 23.44% | 25.94% | 0.75 | 0.73% | 35.56% | 63.72% | 1.63 | 3.36 |
| Shikra (Vicuna-7B) [shikra] | 21.14% | 68.33% | 10.52% | 0.89 | 30.33% | 28.30% | 41.37% | 1.11 | 1.14% | 64.36% | 34.50% | 1.33 | 3.34 |
| Otter-v1 (MPT-7B) [otter] | 22.38% | 59.36% | 18.25% | 0.96 | 40.68% | 35.99% | 23.33% | 0.83 | 1.95% | 13.20% | 84.85% | 1.83 | 3.61 |
| Kosmos-2 [kosmos2] | 8.76% | 70.91% | 20.33% | 1.12 | 29.45% | 34.75% | 35.81% | 1.06 | 0.16% | 14.77% | 85.06% | 1.85 | 4.03 |
| InstructBLIP (Flan-T5-XL) [iblip] | 23.16% | 66.44% | 10.40% | 0.87 | 34.85% | 26.03% | 39.12% | 1.04 | 14.71% | 59.87% | 25.42% | 1.11 | 3.02 |
| InstructBLIP (Vicuna-7B) [iblip] | 29.73% | 61.47% | 8.80% | 0.79 | 27.84% | 23.52% | 48.65% | 1.21 | 27.40% | 61.29% | 11.31% | 0.84 | 2.84 |
| VisualGLM-6B (GLM-6B) [glm] | 30.75% | 56.64% | 12.61% | 0.82 | 38.64% | 26.18% | 35.18% | 0.97 | 6.14% | 67.15% | 26.71% | 1.21 | 2.99 |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | 28.28% | 37.69% | 34.03% | 1.06 | 26.75% | 18.18% | 55.07% | 1.28 | 3.03% | 33.82% | 63.15% | 1.60 | 3.94 |
| LLaMA-Adapter-V2 [llamaadapterv2] | 30.44% | 53.99% | 15.57% | 0.85 | 29.41% | 25.79% | 44.80% | 1.15 | 1.50% | 52.75% | 45.75% | 1.44 | 3.45 |
| LLaVA-v1 (Vicuna-13B) [llava] | 34.10% | 40.52% | 25.39% | 0.91 | 30.02% | 15.15% | 54.83% | 1.25 | 1.06% | 38.03% | 60.91% | 1.60 | 3.76 |
| MiniGPT-4 (Vicuna-13B) [minigpt4] | 34.01% | 32.15% | 33.85% | 1.00 | 29.20% | 15.27% | 55.53% | 1.26 | 6.88% | 45.65% | 47.48% | 1.41 | 3.67 |
@@
| InfiMM (Zephyr-7B) [InfiMM] | 30.75% | 62.66% | 6.22% | 0.75 | 34.17% | 38.84% | 26.35% | 0.91 | 2.57% | 30.84% | 65.28% | 1.61 | 3.28 |
| Emu2-Chat (LLaMA-33B) [emu2] | 41.25% | 54.33% | 4.42% | 0.63 | 38.11% | 36.41% | 25.48% | 0.87 | 4.12% | 38.61% | 57.27% | 1.53 | 3.03 |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 37.95% | 52.17% | 9.11% | 0.70 | 37.68% | 37.33% | 23.73% | 0.84 | 3.95% | 31.15% | 62.84% | 1.56 | 3.12 |
| BakLLava (Mistral-7B) [bakllava] | 29.46% | 59.77% | 10.57% | 0.80 | 40.0% | 38.08% | 21.33% | 0.80 | 2.26% | 15.06% | 82.04% | 1.79 | 3.40 |
| mPLUG-Owl2 (Q-Instruct) [mplug2] | 15.25% | 65.76% | 18.32% | 1.02 | 39.44% | 40.18% | 19.62% | 0.79 | 0.09% | 9.86% | 89.02% | 1.87 | 3.69 |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 19.43% | 65.54% | 14.45% | 0.94 | 30.94% | 43.71% | 24.63% | 0.92 | 3.79% | 26.94% | 68.28% | 1.63 | 3.50 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 19.68% | 72.57% | 7.19% | 0.86 | 38.00% | 40.04% | 20.97% | 0.82 | 2.13% | 39.77% | 56.66% | 1.53 | 3.22 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 18.77% | 73.44% | 7.79% | 0.89 | 34.66% | 38.72% | 26.62% | 0.92 | 1.02% | 34.59% | 64.39% | 1.63 | 3.44 |
@@ III-A2 Perception for Image Pairs
Perception for image pairs is far more difficult for MLLMs since this task not only requires MLLMs to have stable low-level visual capabilities, but also requires MLLMs to be able to analyze two images simultaneously and conduct discerning comparisons. To enrich the MLLM diversity, we further include the mPLUG-Owl2 fine-tuned with the single image low-level visual dataset Q-Instruct [q-instruct] for comparison. The performance is exhibited in Table III. With closer inspections, we can obtain several interesting findings. a) Open-source MLLMs are poor low-level comparators. It seems that although they might show strong performance for single image perception, they are quite confused by the image pairs. Most of them get worse performance on the Compare subset than the Joint subset, which further confirms this point. For mPLUG-Owl2 (Q-Instruct), despite being fine-tuned with the single image low-level visual dataset Q-Instruct [q-instruct], the overall performance improvement from the low-level knowledge infusion of single images is relatively weak. This also suggests that there is a necessity to build open-source low-level datasets for multiple images to cultivate the corresponding capabilities of open-source MLLMs. b) Close-source MLLMs are more robust in this task. This may be because these close-source MLLMs are supported by training on multiple-image data, allowing them to make better comparative judgments. Particularly with GPT-4V, its performance in the compare subset is significantly higher than in the joint subset, and it far exceeds all other models, even reaching the level of a junior human. c) Perception for image pairs is easier for humans. Comparing image pairs is simpler for humans, as the answers to related questions tend to be more objective. Especially for junior-level humans with no professional experience, they may have stronger subjectivity in grasping absolute sensations, but it is easier to remain objective when they are faced with comparison concepts. For example, it’s difficult for a junior-level human to judge whether the lighting in a dimly lit single image is appropriate. However, if presented with another image with even weaker lighting, they can easily determine which image is worse. This may explain the notable 6% improvements for junior-level human from single images to image pairs.
@@ III-B1 Description for single images
| NIQE [niqe] | 0.316/0.377 | 0.693/0.669 | 0.211/0.288 | 0.480/0.451 | 0.075/0.056 | 0.562/0.517 | 0.374/0.428 | 0.387/0.398 |
| CLIP-ViT-Large-14 [clip] | 0.468/0.505 | 0.385/0.389 | 0.218/0.237 | 0.307/0.308 | 0.285/0.290 | 0.436/0.458 | 0.376/0.388 | 0.354/0.368 |
@@
| Emu2-Chat (LLaMA-33B) [emu2] | 0.664/0.714 | 0.712/0.698 | 0.355/0.341 | 0.597/0.611 | 0.224/0.269 | 0.759/0.751 | 0.841/0.790 | 0.593/0.596 |
| Fuyu-8B (Persimmon-8B) [fuyu-8b] | 0.124/0.123 | 0.125/0.179 | 0.164/0.133 | 0.225/0.176 | 0.118/0.116 | 0.368/0.317 | 0.099/0.088 | 0.174/0.161 |
| BakLLava (Mistral-7B) [bakllava] | 0.389/0.390 | 0.406/0.398 | 0.227/0.216 | 0.335/0.337 | 0.179/0.209 | 0.542/0.561 | 0.344/0.361 | 0.346/0.353 |
| mPLUG-Owl2 (LLaMA-7B) [mplug2] | 0.196/0.252 | 0.589/0.614 | 0.217/0.286 | 0.293/0.342 | -0.024/-0.032 | 0.473/0.492 | 0.541/0.546 | 0.326/0.357 |
| LLaVA-v1.5 (Vicuna-v1.5-7B) [improvedllava] | 0.463/0.459 | 0.443/0.467 | 0.305/0.321 | 0.344/0.358 | 0.321/0.333 | 0.672/0.738 | 0.417/0.440 | 0.424/0.445 |
| LLaVA-v1.5 (Vicuna-v1.5-13B) [improvedllava] | 0.448/0.460 | 0.563/0.584 | 0.310/0.339 | 0.445/0.481 | 0.285/0.297 | 0.664/0.754 | 0.390/0.400 | 0.444/0.474 |
| InternLM-XComposer-VL (InternLM) [xcomposer] | 0.564/0.615 | 0.730/0.750 | 0.360/0.416 | 0.612/0.676 | 0.243/0.265 | 0.732/0.775 | 0.546/0.572 | 0.541/0.581 |
| IDEFICS-Instruct (LLaMA-7B) [idefics] | 0.375/0.400 | 0.474/0.484 | 0.235/0.240 | 0.409/0.428 | 0.244/0.227 | 0.562/0.622 | 0.370/0.373 | 0.381/0.396 |
| Qwen-VL (QwenLM) [Qwen-VL] | 0.470/0.546 | 0.676/0.669 | 0.298/0.338 | 0.504/0.532 | 0.273/0.284 | 0.617/0.686 | 0.486/0.486 | 0.475/0.506 |
| Shikra (Vicuna-7B) [shikra] | 0.314/0.307 | 0.320/0.337 | 0.237/0.241 | 0.322/0.336 | 0.198/0.201 | 0.640/0.661 | 0.324/0.332 | 0.336/0.345 |
| Otter-v1 (MPT-7B) [otter] | 0.406/0.406 | 0.436/0.441 | 0.143/0.142 | -0.008/0.018 | 0.254/0.264 | 0.475/0.481 | 0.557/0.577 | 0.323/0.333 |
| Kosmos-2 [kosmos2] | 0.255/0.281 | 0.644/0.641 | 0.196/0.195 | 0.358/0.368 | 0.210/0.225 | 0.489/0.491 | 0.359/0.365 | 0.359/0.367 |
| InstructBLIP (Flan-T5-XL) [iblip] | 0.334/0.362 | 0.582/0.599 | 0.248/0.267 | 0.113/0.113 | 0.167/0.188 | 0.378/0.400 | 0.211/0.179 | 0.290/0.301 |
| InstructBLIP (Vicuna-7B) [iblip] | 0.359/0.437 | 0.683/0.689 | 0.200/0.283 | 0.253/0.367 | 0.263/0.304 | 0.629/0.663 | 0.337/0.382 | 0.389/0.446 |
| VisualGLM-6B (GLM-6B) [glm] | 0.247/0.234 | 0.498/0.507 | 0.146/0.154 | 0.110/0.116 | 0.209/0.183 | 0.342/0.349 | 0.127/0.131 | 0.240/0.239 |
| mPLUG-Owl (LLaMA-7B) [mplugowl] | 0.409/0.427 | 0.634/0.644 | 0.241/0.271 | 0.437/0.487 | 0.148/0.180 | 0.687/0.711 | 0.466/0.486 | 0.432/0.458 |
| LLaMA-Adapter-V2 [llamaadapterv2] | 0.354/0.363 | 0.464/0.506 | 0.275/0.329 | 0.298/0.360 | 0.257/0.271 | 0.604/0.666 | 0.412/0.425 | 0.381/0.417 |
| LLaVA-v1 (Vicuna-13B) [llava] | 0.462/0.457 | 0.442/0.462 | 0.264/0.280 | 0.404/0.417 | 0.208/0.237 | 0.626/0.684 | 0.349/0.372 | 0.394/0.416 |
| MiniGPT-4 (Vicuna-13B) [minigpt4] | 0.239/0.257 | 0.238/0.253 | 0.170/0.183 | 0.339/0.340 | 0.252/0.246 | 0.572/0.591 | 0.239/0.233 | 0.293/0.300 |